The differece between two numbers is 8. Five times the larger number is 4 more than 8 time the other. Find the numbers ?
Answer by scott8148(5891)

let x=smaller, so x+8=larger

"Five times the larger number" __ 5(x+8) __ 5x+40

"is" __ =

"4 more than 8 time the other" __ 8x+4

5x+40=8x+4 __ 36=3x